find K for inconsistent system of equation
4x + 6y = 11
2x + Ky = 7

Answer by ewatrrr(24785)

 
Hi,
Note
A consistent system has at least one solution.
An inconsistent system has no solutions. *****
An independent system has exactly one solution.
A dependent system has infinitely many solutions.
find K for inconsistent system of equation
4x + 6y = 11
2x + Ky = 7 K = 3
4x + 6y = 11
2x + 3y = 7 |multiplying thru by -2 to eliminate the x variable
4x + 6y = 11
-4x +-6y = -14 |multiplying thru by -2 to eliminate the x variable
0 ≠ -4 NO solutions